### v2

# On the Relationship Between Monotone and Squared Probabilistic Circuits

Abstract
Probabilistic circuits are a unifying representation of functions as computation graphs of weighted sums and products. Their primary application is in probabilistic modeling, where circuits with non-negative weights (monotone circuits) can be used to represent and learn density/mass functions, with tractable marginal inference. Recently, it was proposed to instead represent densities as the square of the circuit function (squared circuits); this allows the use of negative weights while retaining tractability, and can be exponentially more expressive efficient than monotone circuits. Unfortunately, we show the reverse also holds, meaning that monotone circuits and squared circuits are incomparable in general. This raises the question of whether we can reconcile, and indeed improve upon the two modeling approaches. We answer in the positive by proposing Inception PCs, a novel type of circuit that naturally encompasses both monotone circuits and squared circuits as special cases, and employs complex parameters. Empirically, we validate that Inception PCs can outperform both monotone and squared circuits on a range of tabular and image datasets.
1 Introduction
The philosophy of tractable probabilistic modeling advocates for modeling high-dimensional probability distributions using model architectures that support exact and efficient inference for various probabilistic queries by-design. This property makes them extremely useful for probabilistic reasoning; for example, tractable models have found wide-ranging applications from enhancing and controlling intractable generative models (Zhang et al. 2023; Liu, Niepert, and Van den Broeck 2024), to neuro-symbolic AI (Ahmed et al. 2022; Ahmed, Chang, and Van den Broeck 2023), to causal discovery and inference (Wang, Wicker, and Kwiatkowska 2022; Wang and Kwiatkowska 2023).
The lingua franca for tractable models is the probabilistic circuits framework (PC) (Choi, Vergari, and Van den Broeck 2020), which specify functions using computation graphs of sums and products, unifying many previous tractable models such as arithmetic circuits (Darwiche 2003), sum-product networks (Poon and Domingos 2011) and cutset networks (Rahman, Kothalkar, and Gogate 2014). The standard approach to modeling is to directly represent the probability distribution (density/mass function) using a PC. In order to enforce non-negativity of the PC output, one typically restricts to the PC to have non-negative parameters; these are known as monotone PCs (Darwiche 2003; Poon and Domingos 2011). However, recent works have also shown that there exist many tractable classes of probability distributions that provably cannot be expressed in this way (Zhang, Holtzen, and Van den Broeck 2020; Yu, Trapp, and Kersting 2023; Broadrick, Zhang, and Van den Broeck 2024).
This motivates the development of new approaches for practically constructing and learning generalized tractable models. To this end, Loconte et al. (2024b) recently proposed squared circuits, where the probability distribution is defined to be (proportional to) the square of the circuit function. In this formulation, a PC can employ real (possibly negative parameters) while still defining a valid probability distribution. This was shown theoretically to lead to an exponential advantage in expressive efficiency compared to monotone PCs, for certain classes of distributions.
In this work, we reexamine monotone and squared (structured-decomposable) PCs, and show that they are incomparable in general: either can be exponentially more expressive efficient than the other. Motivated by this observation, we draw on the latent variable interpretation of PCs (Peharz et al. 2016) and show an elegant connection between the two types of circuits; namely, that they simply correspond to summing the latent variables outside or inside the square, i.e. a deep sum-of-squares or square-of-sums. By combining these two types of latent variables, we introduce a novel class of tractable models representing deep sum-of-square-of-sums, which we call Inception PCs (IncPC), strictly generalizing and extending (structured-decomposable) monotone and squared PCs.
We further investigate learning Inception PCs effectively at scale. Following recent trends in PC learning (Peharz et al. 2020b, a; Liu and Van den Broeck 2021; Mari, Vessio, and Vergari 2023), we design an efficient tensorized implementation of Inception PCs that subsumes tensorized (structured-decomposable) monotone and squared PCs. Empirical results validate the improved expressive efficiency of Inception PCs on density estimation benchmarks.
Our contributions can be summarized as follows:
  • Theoretically, we show that monotone circuits can be exponentially smaller than squared circuits for a simple class of distributions, meaning that monotone and squared circuits are incomparable in general in terms of expressive efficiency (Section 3);
  • We analyze both monotone and squared circuits from a latent variable perspective, showing that they can be interpreted as deep sums-of-squares and squares-of-sums respectively. We then propose a novel class of tractable circuits, Inception PCs (IncPC), which generalize and extend both types of circuits as a deep sum-of-square-of-sums and can employ complex parameters (Section 4).
  • We propose an efficient tensorized architecture for Inception PCs, enabling their application to large-scale datasets (Section 5);
  • Empirically, we demonstrate that (i) combining both types of latents, together with complex parameters, can improve performance; and (ii) Inception PCs can outperform monotone PCs and squared PCs on challenging benchmarks including downscaled versions of ImageNet, even when normalized for computation time (Section 7).
In concurrent work, Loconte, Mengel, and Vergari (2025) proved an exponential separation between monotone and squared circuits (i.e., our Theorem 2) using a similar family of separating functions. They also proposed two new model classes, SOCS and $\mu$SOCS, which generalize squared circuits. We discuss in detail connections with our Inception PCs in Appendix A.
2 Preliminaries
Notation
We use capital letters to denote variables and lowercase to denote their assignments/values (e.g. $X,x$). We use boldface (e.g. $\bm{X},\bm{x}$) to denote sets of variables/assignments.
Probabilistic circuits are computation graphs representing functions constructed by hierarchical compositions of weighted sums and products.
Definition 1 (Probabilistic Circuit).
A probabilistic circuit $\mathcal{C}$ over a set of variables $\bm{V}$ is a rooted DAG consisting of three types of nodes $n$: input, product and sum nodes. Each input node $n$ is a leaf encoding a function $f_{n}:\bm{W}\to\mathbb{R}$ for some $\bm{W}\subseteq\bm{V}$, and for each internal (product or sum) node $n$, denoting the set of children (i.e. nodes $n^{\prime}$ for which $n\rightarrow n^{\prime}$) by $\textnormal{in}(n)$, we define:
$f_{n}=\begin{cases}\prod_{n_{i}\in\textnormal{in}(n)}f_{n}&\text{if }n\text{ is product;}\\ \sum_{n_{i}\in\textnormal{in}(n)}\theta_{n,n_{i}}f_{n_{i}}&\text{if }n\text{ is sum.}\\ \end{cases}$ (1)
where each sum node has a set of weights $\{\theta_{n,n_{i}}\}_{n_{i}\in\textnormal{in}(n)}$ with $\theta_{n,n_{i}}\in\mathbb{R}$. Each node $n$ thus encodes a function over a set of variables $\textnormal{sc}(n)$, which we call its scope; this is given by $\textnormal{sc}(n)=\bigcup_{n_{i}\in\textnormal{in}(n)}\textnormal{sc}(n_{i})$ for internal nodes. The function encoded by the circuit $f_{\mathcal{C}}$ is the function encoded by its root node. The size of a probabilistic circuit $|\mathcal{C}|$ is defined to be the number of edges in its DAG.
In this paper, we will assume that sum and product nodes alternate; this is without loss of generality as this property can be enforced on a PC at most a linear increase in size. A key feature of the sum-product structure of probabilistic circuits is that they allow for efficient (linear-time) computation of marginals, for example the partition function $Z=\sum_{\bm{v}}f(\bm{v})$111alternatively, $\int$ in the case of continuous variables, if they are smooth and decomposable:
Definition 2 (Smoothness, Decomposability).
A probabilistic circuit is smooth if for every sum node $n$, its inputs $n_{i}$ have the same scope. A probabilistic circuit is decomposable if for every product node $n$, its inputs have disjoint scope.
We will also need a stronger version of decomposability that enables circuits to be multiplied together efficiently (Pipatsrisawat and Darwiche 2008; Vergari et al. 2021):
Definition 3 (Structured Decomposability).
A smooth and decomposable probabilistic circuit is structured-decomposable if any two product nodes $n,n^{\prime}$ with the same scope decompose in the same way.
Structured-decomposability is commonly enforced when learning PCs from data (Liu and Van den Broeck 2021) as well as when compiling a circuit from another model such as a Bayesian network (Choi, Kisa, and Darwiche 2013).
3 Expressive Efficiency of Monotone and Squared Structured-Decomposable Circuits
One of the primary applications of probabilistic circuits is as a tractable representation of probability distributions. As such, we typically require the function output of the circuit to be a non-negative real. The usual way to achieve this is to enforce non-negativity of the weights and input functions:
Definition 4 (Monotone PC).
A probabilistic circuit is monotone if all weights are non-negative reals, and all input functions map to the non-negative reals.
Given a monotone PC $\mathcal{C}$, one can define a probability distribution $p_{1}(\bm{V}):=\frac{f_{\mathcal{C}}(\bm{V})}{Z_{\mathcal{C}}}$ where $Z_{\mathcal{C}}$ is the partition function of the PC. However, this is not the only way to construct a non-negative function. In Loconte et al. (2024b), it was proposed to instead use $f_{\mathcal{C}}$ to represent a real (i.e. possibly negative) function, by allowing for real weights/input functions; this can then be squared to obtain a non-negative function. That is, we define $p_{2}(\bm{V}):=\frac{f_{\mathcal{C}}(\bm{V})^{2}}{\sum_{\bm{v}}f_{\mathcal{C}}(\bm{v})^{2}}$.
In order for $\sum_{\bm{v}}f_{\mathcal{C}}(\bm{v})^{2}$ to be tractable to compute, a sufficient condition is for the circuit $\mathcal{C}$ to be structured-decomposable; one can then explicitly construct a smooth and (structured-)decomposable circuit $\mathcal{C}^{2}$ such that $f_{\mathcal{C}^{2}}(\bm{V})=f_{\mathcal{C}}(\bm{V})^{2}$ of size and in time $O(|\mathcal{C}|^{2})$ (Vergari et al. 2021). Then we have that $p_{2}(\bm{V})=\frac{f_{\mathcal{C}^{2}}(\bm{V})}{Z_{\mathcal{C}^{2}}}$, i.e. the distribution induced by the PC $\mathcal{C}^{2}$. Crucially, the circuit $\mathcal{C}^{2}$ is not necessarily monotone; squaring thus provides an alternative means of constructing PCs that represent non-negative functions. In fact, it is known that squared real PCs can be exponentially more expressive efficient than structured-decomposable monotone PCs for representing probability distributions:
Theorem 1.
(Loconte et al. 2024b) There exists a class of non-negative functions $p(\bm{V})$ such that there exist structured-decomposable PCs $\mathcal{C}$ with $p(\bm{V})=f_{\mathcal{C}}(\bm{V})^{2}$ of size polynomial in $|\bm{V}|$, but the smallest structured-decomposable monotone PC $\mathcal{C}^{\prime}$ such that $p(\bm{V})=f_{\mathcal{C}^{\prime}}(\bm{V})$ has size $2^{\Omega(|\bm{V}|)}$.
However, we now show that, in fact, the other direction also holds: monotone PCs can also be exponentially more expressive efficient than squared (real) PCs.
Theorem 2.
There exists a class of non-negative functions $p(\bm{V})$, such that there exist monotone structured-decomposable PCs $\mathcal{C}$ with $p(\bm{V})=f_{\mathcal{C}}(\bm{V})$ of size polynomial in $|\bm{V}|$, but the smallest structured-decomposable PC $\mathcal{C}^{\prime}$ such that $p(\bm{V})=f_{\mathcal{C}^{\prime}}(\bm{V})^{2}$ has size $2^{\Omega(|\bm{V}|)}$.
Proof.
(Sketch) The function class we use to separate the circuit classes is $p(\bm{V})=\sum_{i=0}^{d-1}2^{i}\mathds{1}_{V_{i}=1}+1$, i.e. a function that outputs the integer encoded in binary by the variables ($+1$). This can easily be represented as a structured-decomposable monotone circuit of linear size. We show a lower bound on the size of structured-decomposable circuits computing any square root $F(\bm{V})$ of this function, by (i) reducing to bounding the rank of a matrix representation of $F$ w.r.t. balanced partitions $(\bm{X},\bm{Y})$ of $\bm{V}$ (Martens and Medabalimi 2014), using a standard techinque from communication complexity; (ii) lower bounding the number of distinct prime square roots in the matrix; (iii) showing the existence of a sufficiently large submatrix with full rank, thus lower bounding the rank of the original matrix (Fawzi et al. 2015). ∎
This is perhaps surprising, as squaring PCs generate structured PCs with possibly negative weights, suggesting that they should be more general than monotone structured PCs. The key point is that not all circuits that represent a positive function (not even all monotone structured ones) can be generated by squaring. Taken together, these results are somewhat unsatisfying, as we know that there are some distributions better represented by an unsquared monotone PC, and some by a squared real PC. In the next section, we will investigate how to reconcile these different approaches to specifying probability distributions.
4 Towards a Unified Model for Deep Sums-of-Squares-of-Sums
In this section, we investigate the relationship between monotone and squared circuits in depth. Firstly, we show how to introduce complex parameterizations to squared circuits (Prop 1). Secondly, we provide a new interpretation of monotone and squared circuits as different ways of marginalizing out latent variables (Figure 1); and introduce a new tractable model, Inception PCs, that combines the two approaches (Theorem 3).
4.1 Complex Parameters
We begin by noting that, beyond simply negative parameters, one can also allow for complex weights and input functions [^2] Circuits with complex-valued input functions have previously been proposed by Yu, Trapp, and Kersting (2023) in the context of representing characteristic functions (an alternative representation of probability distributions)., i.e. take values in the field $\mathbb{C}$. Then, to ensure the non-negativity of the squared circuit, we multiply a circuit with its complex conjugate. That is:
$p_{2}(\bm{V})=\frac{|f_{\mathcal{C}}(\bm{V})|^{2}}{\sum_{\bm{v}}|f_{\mathcal{C}}(\bm{v})|^{2}}=\frac{\overline{f_{\mathcal{C}}(\bm{V})}f_{\mathcal{C}}(\bm{V})}{\sum_{\bm{v}}\overline{f_{\mathcal{C}}(\bm{v})}f_{\mathcal{C}}(\bm{v})}$
As complex conjugation is a field isomorphism of $\mathbb{C}$, taking a complex conjugate of a circuit is as straightforward as taking the complex conjugate of each weight and input function, retaining the same DAG as the original circuit. This allows us to efficiently compute $p_{2}(\bm{V})$ as a (smooth and structured decomposable) circuit:
Proposition 1 (Tractability of Complex Conjugation).
Given a smooth and decomposable circuit $\mathcal{C}$, it is possible to compute a smooth and decomposable circuit $\overline{\mathcal{C}}$ such that $f_{\overline{\mathcal{C}}}(\bm{V})=\overline{f_{\mathcal{C}}(\bm{V})}$ of size and in time $O(|\mathcal{C}|)$. Further, if $\mathcal{C}$ is structured decomposable, then it is possible to compute a smooth and structured decomposable $\mathcal{C}^{2}$ s.t. $f_{\mathcal{C}^{2}}(\bm{V})=\overline{f_{\mathcal{C}}(\bm{V})}f_{\mathcal{C}}(\bm{V})$ of size and in time $O(|\mathcal{C}|^{2})$.
4.2 Deep Sums-of-Squares-of-Sums: A Latent Variable Interpretation
$\bm{+}$$\bm{\times}$$\bm{\times}$$\bm{\circ}$$\bm{\circ}$$\llbracket Z=0\rrbracket$$\llbracket Z=1\rrbracket$$3+2i$$1-i$
(a) LVI of sum node $n$
$\bm{+}$$\bm{\times}$$\bm{\times}$131313132222
(b) Computing $\sum_{Z}|f_{n}(\bm{V},Z)|^{2}$
[…]
(c) Computing $|\sum_{Z}f_{n}(\bm{V},Z)|^{2}$
Figure 1: Latent variable interpretation for squaring PCs. The sum node in Figure 1(a) has two children with complex weights and associated with different values of the latent $Z$. A sum-of-squares (Figure 1(b)) gives a monotone PC, where the parameters necessarily become non-negative. A square-of-sums (Figure 1(c)) leads to a squared PC, with four children each corresponding to the product of any two children from the original circuit.
In the latent variable interpretation (LVI) of probabilistic circuits (Peharz et al. 2016), for every sum node, one assigns a categorical latent variable, where each state of the latent variable is associated with one of the inputs to the sum node; we show an example in Figure 1(a). In this interpretation, when performing inference in the probabilistic circuit, we marginalize over all of the latent variables beforehand.
In the case when the circuit is structured decomposable, one can assign latent variables with variable scopes (sets) appearing in the circuit, such that two sum nodes $n,n^{\prime}$ with the same scope $\textnormal{sc}(n)=\textnormal{sc}(n^{\prime})$ are associated with the same latent $Z_{\textnormal{sc}(n)}$. Then, writing $\bm{Z}$ for the set of all latents, the function represented by the PC can be expressed as:
$f_{\mathcal{C}}(\bm{V})=\sum_{\bm{z}}f_{\mathcal{C}}(\bm{V},\bm{z})$ (2)
where $f_{\mathcal{C}}(\bm{V},\bm{z})$ is a product of input functions for any value of $\bm{z}$ [^3] Concretely, $f_{\mathcal{C}}(\bm{V},\bm{z})$ is a function obtained by traversing the circuit top-down, selecting one child of every sum node according to the value $\bm{z}$, and all children of a prouct node; cf. also the notion of induced subcircuits/trees (Chan and Darwiche 2006; Zhao, Poupart, and Gordon 2016). In other words, the distribution represented by the PC is fully factorized conditioned on a latent value $\bm{z}$.
However, interpreting these latent variables becomes tricky when we consider probability distributions defined by squaring circuits. The key question is, does one marginalize out the latent variables before or after squaring? We show both options in Figures 1(b) and 1(c). In Figure 1(b), we square before marginalizing $Z$. In this case, each sum node weight is multiplied by its conjugate, and we are left with a sum node with non-negative real parameters. On the other hand, if we marginalize before squaring, we have a sum node with four children and complex parameters. Interestingly, the former case is very similar to directly constructing a monotone PC, while the latter is more like an explicit squaring without latent variables. This suggests that we can switch between monotone and squared PCs simply by deciding whether to sum the latent variables inside or outside the square. Using this perspective, we propose the following model, which explicitly introduces both type of latent variables into the circuit, producing a deep sum-of-square-of-sums:
Definition 5 (Inception PC [^4] The nomenclature is inspired by the deep(er) layering of summation and squaring in Inception PCs.).
An Inception PC (IncPC) $\mathcal{C}_{\text{Inception}}$ is a smooth and structured-decomposable probabilistic circuit over observed variables $\bm{V}\cup\bm{U}\cup\bm{W}$. The probability distribution of an Inception PC is defined by:
$p_{\text{Inception}}(\bm{V})=\frac{\sum_{\bm{u}}\left|\sum_{\bm{w}}f_{\mathcal{C}_{\text{Inception}}}(\bm{V},\bm{u},\bm{w})\right|^{2}}{\sum_{\bm{v}}\sum_{\bm{u}}\left|\sum_{\bm{w}}f_{\mathcal{C}_{\text{Inception}}}(\bm{v},\bm{u},\bm{w})\right|^{2}}$ (3)
As $\bm{U}$-latents are outside the square, and $\bm{W}$-latents are inside the square, we will refer to them as 1-norm and 2-norm latents respectively. The next Theorem shows that, given an Inception PC, we can efficiently “materialize” it into a tractable PC over just the observed variables $\bm{V}$ representing the Inception PC’s distribution $p_{\text{Inception}}(\bm{V})$:
Theorem 3 (Tractability of InceptionPC).
Given an Inception PC $\mathcal{C}_{\text{Inception}}$, it is possible to compute a smooth and structured decomposable circuit $\mathcal{C}_{\text{mat}}$ such that $f_{\mathcal{C}_{\text{mat}}}(\bm{V})=\sum_{\bm{u}}\left|\sum_{\bm{w}}f_{\mathcal{C}_{\text{Inception}}}(\bm{V},\bm{u},\bm{w})\right|^{2}$ of size and in time $O(|\mathcal{C}|^{2})$.
Proof.
The most systematic way to see this is to take advantage of the compositional inference framework of Vergari et al. (2021). We can marginalize out $\bm{W}$ from $\mathcal{C}_{\text{Inception}}$ to obtain a PC $\mathcal{C}_{1}$ such that $f_{\mathcal{C}_{1}}(\bm{V},\bm{U})=\sum_{\bm{w}}f_{\mathcal{C}_{\text{Inception}}}(\bm{V},\bm{U},\bm{w})$, retaining smoothness and structured decomposability. Then the computation of the square is possible by Proposition 1, returning a smooth and structured-decomposable circuit $\mathcal{C}_{2}$ such that $f_{\mathcal{C}_{2}}(\bm{V},\bm{U})=|\sum_{\bm{w}}f_{\mathcal{C}_{\text{Inception}}}(\bm{V},\bm{U},\bm{w})|^{2}$. Finally, we can marginalize out $\bm{U}$ from this circuit to obtain the structured decomposable and smooth circuit $\mathcal{C}_{\text{mat}}$. ∎
In Figure 2, we show a (fragment of) an example Inception PC, and its materialization via Theorem 3. In the rest of the paper, we will refer to this as the materialized Inception PC. The point is that we can efficiently perform inference on the distribution of an Inception PC through standard PC inference procedures on the materialized IncPC. Though this is not strictly necessary, we will assume (in accordance with the LVI for monotone and squared PCs) that $\bm{W},\bm{U}$ are categoricals, and the input nodes with scope over these variables are indicators (e.g. $\llbracket W=1\rrbracket$).
This provides an elegant resolution to the tension between monotone and squared (real/complex) PCs. To retrieve a monotone PC, we need only set $\bm{W}=\emptyset$; then there is no summation inside the square, and $\mathcal{C}_{\text{mat}}$ has the same structure as $\mathcal{C}_{\text{Inception}}$ but with the parameters and input functions squared (and so non-negative real). [^5] Interestingly enough, in this case we can relax the conditions of Theorem 3 to require decomposability rather than structured decomposability, assuming that children of sum nodes correspond to different indicators and thus $\mathcal{C}_{\text{Inception}}$ is deterministic. Multiplying a deterministic circuit with itself (or its conjugate) is tractable in linear time (Vergari et al. 2021). To retrieve a squared PC, we simply set $\bm{U}=\emptyset$; then there is no summation outside the square. However, by using both types of latents, we obtain a generalized PC model that is strictly more expressive efficient than either individually [^6] The special cases can also be learned in the general case by making $f_{\mathcal{C}_{\text{Inception{}}}}$ constant w.r.t. $\bm{U}$ or $\bm{W}$..
[…]
(a) Inception PC $\mathcal{C}_{\text{Inception{}}}$
[…]
(b) Materialized Inception PC $\mathcal{C}_{\text{mat}}$
Figure 2: Diagrams showing an Inception PC, and the corresponding materialized IncPC. Each product node is labelled with an index, such that e.g. ×$\times_{34}$ is the product of the product nodes ×$\times_{3},\times_{4}$. For clarity, the children of product nodes have been omitted (except the latent indicators), and edge weights for the materialized Inception PC are displayed below the corresponding child.
Corollary 1 (Expressive Efficiency of Inception PCs).
Inception PCs are strictly more expressive efficient than both (structured-decomposable) monotone and squared PCs.
5 Tensorized Inception PCs
Thus far, we have described a general formulation of Inception PCs that simply requires smoothness and structured decomposability of the circuit. In practice, we will want to design specific circuit architectures for learning. The purpose of this section is to (i) propose such an architecture suitable for learning and inference on GPUs; and (ii) present an alternative view of Inception PCs as performing hierarchical tensor contractions.
We follow recent trends in probabilistic circuit learning (Peharz et al. 2020a; Mari, Vessio, and Vergari 2023) and consider tensorized architectures for $\mathcal{C}_{\text{Inception{}}}$, where sum and product nodes are grouped into regions by scope. In particular, for each region (scope), we construct $N_{1}\times N_{2}$ sum and product nodes in the Inception PC, where the sum nodes are connected in a dense fashion to the product nodes via a weight tensor $\theta_{ii^{\prime}kk^{\prime}}\in\mathbb{C}^{N_{1}^{2}\times N_{2}^{2}}$ with $1\leq k,k^{\prime}\leq N_{1}$, $1\leq i,i^{\prime}\leq N_{2}$; and the product nodes are connected to sum nodes in subsequent regions via a Hadamard product (Loconte et al. 2024a). Each region is associated with a categorical latent $W$ and $U$; and each product node corresponds explicitly to a value of $W$ and $U$, having two indicators $\llbracket W=i^{\prime}\rrbracket$ and $\llbracket U=k^{\prime}\rrbracket$ as children where $1\leq k^{\prime}\leq N_{1}$, $1\leq i^{\prime}\leq N_{2}$. For example, the product nodes in the Inception PC in Figure 2(a) correspond to a tensorized Inception PC with $N_{1}=2,N_{2}=2$.
Once materialized using Theorem 3, we obtain a materialized IncPC $\mathcal{C}_{\text{mat}}$ with $N_{1}\times N_{2}^{2}$ sum and product nodes; the increase in size occurs from the “expansion” from two-norm latents $\bm{W}$ as seen in Figure 1(c). For a given region, let us write $n^{(S)}_{ijk}$ for the sum nodes for $1\leq k\leq N_{1},1\leq i,j\leq N_{2}$, and $n^{(P)}_{ijk}$ for the product nodes for $1\leq k\leq N_{1},1\leq i,j\leq N_{2}$ in $\mathcal{C}_{\text{mat}}$. The product nodes are still connected to their $M$ child sum regions $n^{(S_{1})},...,n^{(S_{M})}$ via a Hadamard product:
$f_{n^{(P)}_{ijk}}=\prod_{m=1}^{M}f_{n^{(S_{m})}_{ijk}}$ (4)
However, the sum-to-product connection is no longer dense. In particular, writing we have the following relationship between the sum and product node tensors:
$f_{n^{(S)}_{ijk}}=\sum_{i^{\prime}j^{\prime}k^{\prime}}\theta_{ii^{\prime}kk^{\prime}}\overline{\theta_{jj^{\prime}kk^{\prime}}}f_{n^{(P)}_{i^{\prime}j^{\prime}k^{\prime}}}$ (5)
We illustrate the structure of this sum region in Figure 3. On a high level, the connections between the groups of nodes in Figure 3 can be viewed as a standard, monotone PC; in particular, the value of each sum group is simply an weighted sum of its children. However, each weighted edge is between 2D groups of “squared” nodes, rather than between scalar nodes.
$\bm{+}$$\bm{+}$$\bm{+}$$\bm{+}$$\bm{+}$$\bm{+}$$\bm{+}$$\bm{+}$$\bm{+}$$\bm{+}$$\bm{+}$$\bm{+}$$\bm{\times}$$\bm{\times}$$\bm{\times}$$\bm{\times}$$\bm{\times}$$\bm{\times}$$\bm{\times}$$\bm{\times}$$\bm{\times}$$\bm{\times}$$\bm{\times}$$\bm{\times}$$i=1,2$$j=1,2$$k=1$$k=2$$k=3$
Figure 3: Illustration of tensorized Inception PC sum region, where $N_{1}=3$, $N_{2}=2$.
The complexity of a forward pass can be deduced from the tensor contraction in Equation 5. In Table 1, we compare the parameter counts (i.e., memory cost) and time complexity of a forward pass for each region, between monotone, squared and Inception PCs; where $B$ is the size of the data batch. It can be seen that Inception PCs have the same number of parameters/complexity as monotone and squared PCs when setting $N_{2}=1$ and $N_{1}=1$ respectively, except with the complexity differing for squared PCs: this is because one can perform the squaring only once-per-batch in this special case (Loconte et al. 2024b).
As with previous works (Peharz et al. 2020a), we organize the circuit into layers (sets of regions that can be computed simultaneously) to further take advantage of GPU parallelization. For training, we use gradient descent on the negative log-likelihood of the training set. In the case of using complex parameters, we use Wirtinger derivatives (Kreutz-Delgado 2009) in order to optimize the complex weights and input functions. To achieve numerical stability, we use a variant of the log-sum-exp trick for complex numbers, which we describe in Appendix D.
Monotone Squared Inception
Complexity $BN_{1}^{2}$ $BN_{2}^{2}+N_{2}^{3}$ $BN_{1}^{2}N_{2}^{3}$
Parameter Count $N_{1}^{2}$ $N_{2}^{2}$ $N_{1}^{2}N_{2}^{2}$
Table 1: Complexity of batched forward pass and parameter count for circuit variants, per region.
Dataset Model
Monotone PC Squared PC Inception PC RAT-SPN VAE
Non-negative Real Complex Non-negative Real Complex
nltcs 6.04 6.02 6.09 6.03 6.00 6.01 6.02 6.01 5.99
msnbc 6.25 6.23 6.10 6.07 6.05 6.06 6.04 6.04 6.09
kdd-2k 2.13 2.10 2.23 2.12 2.12 2.13 2.12 2.13 2.12
plants 13.70 13.38 15.06 13.13 12.81 13.06 12.76 13.44 12.34
jester 52.77 52.56 54.51 52.97 52.51 52.60 52.75 52.97 51.54
audio 40.27 40.08 41.49 40.01 39.88 39.91 40.05 39.96 38.67
netflix 57.09 56.85 57.68 56.70 56.52 56.57 56.74 56.85 54.73
accidents 29.57 27.93 28.15 27.05 26.70 27.30 26.61 35.49 29.11
retail 10.99 10.82 11.00 10.95 11.00 10.95 10.95 10.91 10.83
pumsb-star 27.98 24.95 25.69 23.98 23.69 24.85 23.03 32.53 25.16
dna 80.21 79.95 80.15 80.17 79.85 80.11 79.77 97.23 94.56
kosarek 10.77 10.54 12.03 10.59 10.69 10.83 10.60 10.89 10.64
msweb 10.44 9.92 10.58 10.17 10.84 10.34 10.10 10.12 9.727
book 33.70 33.32 37.02 33.95 33.51 34.18 33.67 34.68 33.19
eachmovie 52.83 51.28 62.03 52.33 50.76 51.22 51.41 53.63 47.43
web-kb 155.34 151.84 162.03 155.00 151.74 153.32 153.98 157.53 146.9
reuters-52 95.22 92.63 96.25 93.90 93.17 89.67 93.80 87.37 81.33
20ng 155.05 152.98 164.19 154.79 154.15 155.47 155.18 152.06 146.9
bbc 253.98 250.88 259.04 255.13 251.28 253.28 253.37 252.14 240.94
ad 16.93 15.54 16.32 15.35 16.02 15.81 15.32 48.47 18.81
Table 2: Test negative log-likelihoods on 20 binary datasets. Lower is better.
6 Related Work
The space of probabilistic models can be usefully characterized through the lens of tractability and expressive efficiency (Choi, Vergari, and Van den Broeck 2020). Tractability forms a spectrum, extending from intractable models such as diffusion models (Ho, Jain, and Abbeel 2020) and variational autoencoders (Kingma and Welling 2014), to models admitting tractable likelihoods such as normalizing flows (Papamakarios et al. 2021), to models admitting tractable marginals and more complex queries (Vergari et al. 2021; Wang et al. 2024) such as probabilistic circuits. Probabilistic circuits, however, are generally less expressive efficient in practice than less tractable models due to the structure that is imposed in the copmutation graph. Thus much effort has been expended on improving the expressive efficiency of probabilistic circuits whilst maintaining tractability (Sidheekh and Natarajan 2024).
Our work builds upon a long line of work in the circuit literature examining the effect of relaxing the monotonicity condition in circuits. Theoretically, it is known that allowing negative parameters in arithmetic circuits can result in exponential gains in succinctness (Valiant 1979), though this not true of all circuit subclasses (de Colnet and Mengel 2021). Practically, recent works have aimed to exploit negative parameters in probabilistic modeling (Zhang, Holtzen, and Van den Broeck 2020; Zhang, Juba, and Van den Broeck 2021; Sladek, Trapp, and Solin 2023; Loconte et al. 2024b). Yu, Trapp, and Kersting (2023) design circuits to represent the characteristic function of a distribution, which can take complex values (in particular, using complex leaves). Concurrently with our work, Loconte, Mengel, and Vergari (2025) proposed to employ a sum of squared circuits to overcome similar limitations to those we observe in this paper.
There also exist a range of other probabilistic models which employ negative parameterizations and/or squaring. Tensor networks, such as the popular matrix-product states (MPS) (Perez-Garcia et al. 2007), compactly encode functions through sparse tensor contractions. They are often used to model quantum states, whereby the probability of observations is given by squaring via the Born rule (Dirac 1981); recently, tensor networks have also been used for probabilistic modeling (Cheng et al. 2019; Glasser et al. 2019). Positive semidefinite models (Rudi and Ciliberto 2021) in the kernel methods literature utilize a shallow sum of squares to define unnormalized distributions. Tsuchida, Ong, and Sejdinovic (2023, 2024) recently introduced squared neural families, which employ the squared 2-norm of a neural network in the density function and strictly generalize exponential familiies. We discuss technical connections with these models in Appendix A.
7 Experiments
In our experiments, we aim to answer the following research questions: (1) do Inception PCs improve upon the expressivity and modeling capabilities of monotone and squared PCs?; (2) what is the tradeoff between the number of one-norm latents and two-norm latents in terms of modeling performance and computational cost?; (3) how do complex parametererizations of Inception PCs compare to non-negative and real parameterizations?
We use hidden Chow-Liu trees (HCLT) in all experiments as the PC vtree (Liu and Van den Broeck 2021), as it satisfies the required structured-decomposability property, and has been shown to provide state-of-the-art likelihoods for PCs. Further experimental details can be found in Appendix E.
7.1 Binary Datasets
We begin by evaluating on the 20 binary datasets, which has been extensively used as a density estimation benchmark (Rooshenas and Lowd 2014; Peharz et al. 2020b). We aim to investigate across this range of datasets how (1) non-negative, real, and complex parameterizations differ; and (2) whether Inception PCs can effectively make use of both types of latents, by comparing to its monotone and squared PC counterparts. In particular, we set $(N_{1},N_{2})=(8,1)$ for monotone PCs, $(N_{1},N_{2})=(1,8)$ for squared PCs, and $(N_{1},N_{2})=(8,8)$ for Inception PCs. We train each PC on negative log-likelihood for 100 epochs, using the Adam optimizer (Kingma and Ba 2015) with learning rate $0.01$. We average over 5 runs for each configuration.
The results are shown in Table 2. It can be seen that Inception PCs achieve the best likelihoods on 14 of the 20 datasets, confirming that they provide more modeling capacity compared to squaring alone. They also perform very favorably compared to existing models, including (tractable) RAT-SPNs (Peharz et al. 2020b) and (intractable) importance weighed autoencoders (Burda, Grosse, and Salakhutdinov 2016). Squared PCs, even with non-negative parameters, outperform their monotone counterparts that have the same number of parameters, as also observed in Figure 3 of Loconte et al. (2024b). Very interestingly, despite the fact that complex parameterizations of edge weights strictly generalizes real parameterizations and non-negative parameterizations, there is no clear trend in performance, with the non-negative parameterization often achieving comparable or better likelihoods than real or complex parameterizations. We observed that this is (at least partially) due to the complex and negative parameterizations being more prone to overfitting (cf. training LLs in Appendix F).
7.2 Scaling to Large Image Datasets
In this section, we conduct experiments on large-scale image datasets, in particular, downscaled versions of ImageNet to $32\times 32$ and $64\times 64$ (Deng et al. 2009). Following recent work on PC modeling for these datasets (Liu, Zhang, and Van den Broeck 2023; Liu et al. 2023; Liu, Ahmed, and Van den Broeck 2024), we transform the data from RGB using the lossy YCoCg transform. Note that likelihoods on on YCoCg transformed data are thus not comparable to likelihoods on the original RGB dataset. Additionally, to improve training efficiency, we train the circuits on 16 $\times$ 16 patches of the original image [^7] The entire image is then modelled using the same PC for each of the patches; the performance could potentially be improved further by modelling correlation using a PC over patch-level PCs (Liu et al. 2023)., such that the PC models $16\times 16\times 3=768$ variables each with $256$ categories. We train the models on negative log-likelihood, using the Adam optimizer with learning rate $0.01$. We evaluate models using test-set bits-per-dimension (bpd).
We begin by examining the tradeoff between the two types of latents. To this end, in Figure 4 we plot the test bpd for a range of configurations of $(N_{1},N_{2})$ on ImageNet32. These configurations were chosen based on a search in the range $N_{1},N_{2}\in\{1,2,4,8,16,32,64\}$ according to a maximum budget of $2621442^{18}=$ 262144 floating-point operations per second (FLOPS) per region (c.f. Table 1). It can be seen that the optimal configuration ($N_{1}=32,N_{2}=4$ with bpd $4.19$) given these constraints uses a combination of 1-norm latents and 2-norm latents, as opposed to a pure monotone ($N_{2}=1$) or pure squared circuit ($N_{1}=1$).
In Table 3, we summarize results from the ImageNet32 and ImageNet64 datasets, where MPC refers to a monotonic PC $(N_{1},N_{2})=(64,1)$, SQC a squared PC $(N_{1},N_{2})=(1,64)$, and IncPC a tensorized Inception PC with the optimal configuration of $(N_{1},N_{2})$ under the computation time constraint. For comparison, we also show results for monotone HCLT PCs trained using expectation-maximization (EM), and using latent variable distillation (LVD) in which guidance for the PC latent space is provided by distilling information from existing deep generative models (Liu, Zhang, and Van den Broeck 2023). The results show that it is possible to effectively train tractable PC models from initialization with Adam with bpds competitive with the state-of-the-art.
Figure 4: Test bpd for a range of configurations of $(N_{1},N_{2})$ on ImageNet32 (lower is better); configurations are limited to $2^{18}=262K$ FLOPS per region. The Inception PC with $N_{1}=32,N_{2}=4$ achieves the best performance.
MPC SPC IncPC EM LVD
ImageNet32 4.35 4.26 4.19 4.82 4.38
ImageNet64 4.12 3.98 3.90 4.67 4.12
Table 3: Test bpd on large-scale image datasets (lower is better). All PCs use the HCLT vtree.
8 Conclusion
To conclude, we have shown that two important classes of tractable probabilistic models, namely monotone and squared real structured-decomposable PCs are incomparable in terms of expressive efficiency in general. Thus, we propose a new class of probabilistic circuits, Inception PCs, based on deep sums-of-squares-of-sums that generalizes these approaches, and can employ complex parameters. We further show empirically that Inception PCs can offer better performance on a range of tabular and image datasets. Promising avenues to investigate in future work would be improving the optimization of complex parameters in Inception PCs; as well as reducing the computational cost of training by designing more efficient architectures.
Acknowledgements
BW and GVdB gratefully acknowledge support from the National Artificial Intelligence Research Resource Pilot under project NAIRR240143. This work was funded in part by the DARPA ANSR program under award FA8750-23-2-0004, the DARPA PTG Program under award HR00112220005, and NSF grant #IIS-1943641.
Appendix A Relationship with Other Models
In this section, we explain the relationship of Inception PCs with other tractable model classes (i.e. probabilistic models admitting efficient computation of normalizing constants) employing similar principles.
A.1 Probabilistic Circuits
Loconte, Mengel, and Vergari (2025) concurrently introduced sum of compatible squares (SOCS), which models a distribution as a shallow sum of squared circuits:
$p_{\text{SOCS}}(\bm{V})\propto\sum_{i=0}^{n-1}|f_{i}(\bm{V})|^{2}$ (6)
where $f_{1},...,f_{n}$ are probabilistic circuits that are compatible, i.e. share the same scope decomposition. Using the latent variable interpretation $f_{i}(\bm{V})=\sum_{\bm{w}}f_{i}(\bm{V},\bm{w})$, we can interpret this as a special case of an InceptionPC where there is only one 1-norm latent variable $U_{\bm{V}}$ taking values in $10,\ldots,n-$ 1, and with augmented PC function given by
$f_{\mathcal{C}_{\text{aug}}}(\bm{V},u_{\bm{V}},\bm{w})=f_{u_{\bm{V}}}(\bm{V},\bm{w})$ (7)
In addition, the authors also proposed taking the product of compatible monotone and squared circuits ($\mu$SOCS), i.e.
$p_{\mu\text{SOCS}}(\bm{V})\propto f_{1}(\bm{V})|f_{2}(\bm{V})|^{2}$ (8)
Once again, we can employ the latent variable interpretation $f_{1}(\bm{V})=\sum_{u}f_{1}(\bm{V},\bm{u})$ and $f_{2}(\bm{V})=\sum_{\bm{w}}f_{2}(\bm{V},\bm{w})$, which can be represented using the following augmented PC function:
$f_{\mathcal{C}_{\text{aug}}}(\bm{V},\bm{u},\bm{w})$ $=\sqrt{f_{1}(\bm{V},\bm{u})}f_{2}(\bm{V},\bm{w})$ (9)
That is $\mu$SOCS circuits correspond to a particular factorization of the augmented PC over $\bm{u},\bm{w}$, such that the 1-norm and 2-norm latents are enforced to be “independent”. In other words, both SOCS and $\mu$SOCS can be viewed as instances of Inception PCs.
A key advantage compared to general Inception PCs, however, is that they enable the computation of unnormalized likelihoods without explicitly performing the squaring operation. In particular, the log-RHS of equations 6 and 8 can be computed by computing the $f_{i}$ (or $f_{1},f_{2}$) on a given sample $\bm{V}=\bm{v}$ and then computing the expression (similar trick to squared PCs). This is not possible for general Inception PCs, because of the dependence between $\bm{U}$ and $\bm{W}$. Thus, to train e.g. $\mu$SOCS one needs to perform the squaring explicitly only once per batch/parameter update to obtain the normalizing constant. This leads to a training time complexity of $O\bigl{(}N_{1}^{2}N_{2}^{3}+B(N_{1}^{2}+N_{2}^{2})\bigr{)}$ per region, which can be significantly cheaper than the $O(BN_{1}^{2}N_{2}^{3})$ for large batch sizes. For inference queries such as probability or marginal computation, the complexity is the same unless the queries can be batched effectively. The improved time complexity of $\mu$SOCS, however, may come at the cost of expressivity and modeling performance. An interesting open question is thus to determine whether there is a separation between $\mu$SOCS and the general class of all Inception PCs.
A.2 Tensor Networks
In the tensor network literature, it has been proposed both to use non-negative tensor factorizations and the square of complex tensor factorizations (Han et al. 2018) as parameterizations of a tensor with non-negative entries and thus to represent multidimensional discrete probability distributions. While tensor networks are intractable in general, the widely used matrix-product states (MPS) / tensor-train (TT) and tree tensor networks (TTN) admit tractable inference, and can be interpreted as structured-decomposable circuits; in particular, right-linear vtrees for MPS and general vtrees for TTN (Loconte et al. 2024b, a).
Particularly relevant in our context are the locally purified states (LPS) introduced by Glasser et al. (2019), which are a generalization of squared MPS. In the original notation, this factorizes a tensor $T_{X_{1},\ldots,X_{n}}$ over variables (indices) $X_{1},\ldots,X_{n}$ as follows:
$T_{X_{1},...,X_{n}}=\sum_{\{\alpha_{i},\alpha^{\prime}_{i}=1\}}^{r}\sum_{\{\beta_{i}=1\}}^{\mu}A_{1,X_{1}}^{\beta_{1},\alpha_{1}}\overline{A_{1,X_{1}}^{\beta_{1},\alpha_{1}^{\prime}}}A_{2,X_{2}}^{\beta_{2},\alpha_{1},\alpha_{2}}\overline{A_{2,X_{2}}^{\beta_{2},\alpha_{1}^{\prime},\alpha_{2}^{\prime}}}\ldots A_{n,X_{n}}^{\beta_{n},\alpha_{n}}\overline{A_{n,X_{n}}^{\beta_{n},\alpha_{n}^{\prime}}}$ (10)
where $A_{1},...,A_{n}$ are tensors. In contrast, when employing a right-linear vtree, our Inception PCs can be written using the following tensor factorization (matching notation to the extent possible):
$T_{X_{1},...,X_{n}}=\sum_{\{\alpha_{i},\alpha^{\prime}_{i}=1\}}^{r}\sum_{\{\beta_{i}=1\}}^{\mu}A_{1}^{\beta_{1},\alpha_{1}}\overline{A_{1}^{\beta_{1},\alpha_{1}^{\prime}}}E_{1,X_{1}}^{\beta_{1},\alpha_{1}}\overline{E_{1,X_{1}}^{\beta_{1},\alpha_{1}^{\prime}}}A_{2}^{\beta_{1},\beta_{2},\alpha_{1},\alpha_{2}}\overline{A_{2}^{\beta_{1},\beta_{2},\alpha_{1}^{\prime},\alpha_{2}^{\prime}}}E_{2,X_{2}}^{\beta_{2},\alpha_{2}}\overline{E_{2,X_{2}}^{\beta_{2},\alpha^{\prime}_{2}}}\ldots$ (11)
Here, the $A$ tensors correspond to sum node weight matrices, and $E$ tensors correspond to the input categorical functions. Intuitively, one can view the $\alpha_{i}$ indices as corresponding to the 2-norm latents $\bm{W}$ and $\beta_{i}$ to the 1-norm latents $\bm{U}$. Thus Inception PCs share similarities with LPS in extending beyond pure monotone or squared tensor factorizations. The key differences with LPS are that (i) $X_{i}$ only depends on $\alpha_{i},\alpha_{i}^{\prime}$ (not $\alpha_{i-1}$) and (ii) $\beta_{i-1},\beta_{i}$ appear together in the transition tensor.
Thus, with this interpretation, LPS and Inception PCs can be understood essentially as different patterns for constructing a non-negative tensor/probability distribution using complex parameters. LPS could also be in theory generalized to arbitrary vtrees (i.e. tree tensor network) with some work. The practical choice of which to use depends on two main factors. Firstly, it can be checked that the complexity of evaluting the LPS is $O(BN_{1}N_{2}^{3}d)$ per region, where $\mu=N_{1}$, $r=N_{2}$, and $d$ is the dimension of the indices (i.e. the number of categories of each $X$ variable). This is as compared to the $O(BN_{1}^{2}N_{2}^{3})$ complexity for Inception PCs. Secondly, one should consider the performance scaling with the $N_{1}$ and $N_{2}$ parameters. The experiments of Glasser et al. (2019) showed little to no improvement of LPS over Born machines in terms of log-likelihood, which correspond to setting $N_{1}>1$ or $N_{1}=1$ respectively. In contrast, our experiments show significant improvements in log-likelihood when increasing $N_{1}$. This suggests that the Inception PC pattern is more effective with scaling w.r.t. the 1-norm latents.
A.3 Squared Neural Families and PSD models
Positive semi-definite (PSD) kernel models (Marteau-Ferey, Bach, and Rudi 2020; Rudi and Ciliberto 2021) specify probability distributions as:
$p(\bm{x})\propto\bm{\kappa}(\bm{x})^{T}\bm{A}\bm{\kappa}(\bm{x})$ (12)
where $\bm{\kappa}$ is some feature map (which is assumed to be finite-dimensional) and $\bm{A}$ a positive semi-definite matrix. Since this model is shallow, Sladek, Trapp, and Solin (2023) proposed to embed these models within a deep probabilistic circuit architecture by introducing PSD nodes, where the feature map is given by a vector of circuit nodes. That is, the function given by the PSD node is defined as $\bm{f}(\bm{x})^{T}\bm{A}\bm{f}(\bm{x})$ where $\bm{f}(\bm{x})=[f_{1}(\bm{x}),...f_{n}(\bm{x})]$ is a vector of node functions. They considered in particular settings where the PSD nodes are either located at the leaves of the PC, or at the root of the PC. In both such cases one can encode this model efficiently as a $\mu$SOCS and thus an Inception PC (Loconte, Mengel, and Vergari 2025).
Tsuchida, Ong, and Sejdinovic (2023) recently introduced squared neural families (SNEFY), a class of probabilistic models that specify unnormalized probability densities by squaring a neural network’s output, i.e.,
$P(d\bm{x};\bm{V},\bm{\Theta})=\frac{\mu(d\bm{x})}{Z(\bm{V},\bm{\Theta})}||f(\bm{t}(\bm{x});\bm{V},\bm{\Theta})||^{2}$ (13)
where $\bm{X}$ are the variables and $\bm{V}$ and $\bm{\Theta}$ are the neural network parameters, $\bm{t}$ is a sufficient statistic, and $Z(\bm{V},\bm{\Theta})$ is a normalizing constant. Under certain conditions on $f$, $\bm{t}$ and base measure $\mu$, including but not limited to exponential families, it is possible to efficiently compute the normalizing constant of this distribution efficiently. It was shown that some of these SNEFYs can be encoded as $\mu$SOCS (Loconte, Mengel, and Vergari 2025), and thus Inception PCs by Section A.1. Another point of interest is that tractable SNEFYs can be used as expressive conditional probability distributions (i.e. $p(\bm{x}|\bm{y})$) by making the parameters a function of the condition $\bm{y}$ in a particular way. One can also use Inception PCs as a conditional probability model, either by (i) explicitly introducing evidence on a circuit respecting the joint distribution $p(\bm{x},\bm{y})$ – in which case it is not too hard to see that the result remains an Inception PC – or (ii) by similarly making the InceptionPC parameters a (neural network) function of the condition (Shao et al. 2022).
We also refer readers to the excellent technical comparison in (Loconte, Mengel, and Vergari 2025) for further details on the relationship between these model types and circuits.
Appendix B Proofs
See 1
Proof.
We show the first part inductively from leaves to the root. By assumption, we can compute the complex conjugate of the input functions. Thus we need to show that we can compute the conjugate of the sums and products efficiently, assuming that we can compute the conjugates of their inputs.
Suppose that we have a sum $n$; then we have that: $\overline{f_{n}}=\overline{\sum_{n_{i}\in\textnormal{in}(n)}\theta_{n,n_{i}}f_{n_{i}}}=\sum_{n_{i}\in\textnormal{in}(n)}\overline{\theta_{n,n_{i}}}\;\overline{f_{n_{i}}}$. Thus we can simply conjugate the weights and take the conjugated input nodes.
Suppose that we are given a product $n$; then we have that: $\overline{f_{n}}=\overline{\prod_{n_{i}\in\textnormal{in}(n)}f_{n_{i}}}=\prod_{n_{i}\in\textnormal{in}(n)}\overline{f_{n_{i}}}$. Thus we can take the conjugated input nodes.
This procedure is clearly linear time and keeps exactly the same structure as the original circuit (thus smoothness and decomposability). If the input circuit is structured decomposable, then we can multiply $f_{\mathcal{C}}$ and $f_{\overline{\mathcal{C}}}$ as they are compatible (Vergari et al. 2021), producing a smooth and structured decomposable circuit as output. ∎
See 2
Proof.
Given a set of $d$ variables $\bm{V}$, we consider the function:
$p(\bm{V})=n(\bm{V})+1$ (14)
where we write $n(\bm{V})$ for the non-negative integers given by the binary representation.
Existence of Compact Str.Dec.Monotone Circuit
This function can be easily represented as a linear-size monotone structured-decomposable PC as follows:
$p(\bm{V})=f_{\mathcal{C}}(\bm{V})=\sum_{i=0}^{d-1}2^{i}\mathds{1}_{V_{i}=1}+1$
which can also be easily smoothed if desired.
Lower Bound Strategy
It remains to show the lower bound on the size of the negative structured-decomposable PC $\mathcal{C}^{\prime}$. Firstly, we have the following Lemma:
Lemma 1.
(Martens and Medabalimi 2014) Let $F$ be a function over variables $\bm{V}$ computed by a structured-decomposable and smooth circuit $\mathcal{C}$. Then there exists a partition of the variables $(\bm{X},\bm{Y})$ with $\frac{1}{3}|\bm{V}|\leq|\bm{X}|,|\bm{Y}|\leq\frac{2}{3}|\bm{V}|$ and $N\leq|\mathcal{C}|^{2}$ such that:
$F(\bm{X},\bm{Y})=\sum_{i=1}^{N}G_{i}(\bm{X})\times H_{i}(\bm{Y})$ (15)
for some functions $G_{i},H_{i}$.
Variations of this result have appeared multiple times in literature, e.g. Theorem 38 in (Martens and Medabalimi 2014) and Theorem 2 in (de Colnet and Mengel 2021). Unlike prior results, we do not place the restriction that $F$ is non-negative; this is a simple modification of the existing proof as this was only required to ensure the non-negativity of $G_{i},H_{i}$, which we do not require.
To show a lower bound on $|\mathcal{C}^{\prime}|$, we can thus show a lower bound on $N$ for all functions $F(\bm{X},\bm{Y})=\pm\sqrt{n(\bm{V})+1}$. To do this, we use another Lemma (Lemma 13 in de Colnet and Mengel (2021)):
Definition 6.
Given a function $F$ over variables $\bm{X},\bm{Y}$, we define the value matrix $M_{F(\bm{X},\bm{Y})}\in\mathbb{R}^{2^{|\bm{X}|}\times 2^{|\bm{Y}|}}$ by:
$M_{n(\bm{X}),n(\bm{Y})}:=F(\bm{X},\bm{Y})$ (16)
Lemma 2.
(de Colnet and Mengel 2021) Suppose Equation 15 holds. Then $rank(M_{F(\bm{X},\bm{Y})})\leq N$.
Thus, it suffices to lower bound $rank(M_{F(\bm{X},\bm{Y})})$ over all partitions $\bm{X},\bm{Y}$ such that $\frac{1}{3}|\bm{V}|\leq|\bm{X}|,|\bm{Y}|\leq\frac{2}{3}|\bm{V}|$.
Lower Bound
Given such a partition $\bm{X},\bm{Y}$, assume w.l.o.g. $|\bm{X}|\leq|\bm{Y}|$. Consider any function $F(\bm{V})$ such that $F(\bm{V})=\pm\sqrt{n(\bm{V})+1}$.
Each variable $X\in\bm{X}$ corresponds to some variable in $\bm{V}$. We write $idx(X)$ to denote the index of the variable $X$ corresponds to; for example, if $X$ is $V_{4}$, then $idx(X)=4$. Then we have the following:
$F(\bm{X},\bm{Y})=\pm\sqrt{\sum_{i=0}^{|\bm{X}|-1}2^{idx(X_{i})}X_{i}+\sum_{i=0}^{d-|\bm{X}|-1}2^{idx(Y_{i})}Y_{i}+1}$ (17)
We write $\iota(\bm{X}):=\sum_{i=0}^{|\bm{X}|-1}2^{idx(i)}X_{i}$ and $\iota(\bm{Y}):=\sum_{i=0}^{d-|\bm{X}|-1}2^{idx(i)}Y_{i}$ such that $F(\bm{X},\bm{Y})=\pm\sqrt{\iota(\bm{X})+\iota(\bm{Y})+1}$. Note that $\iota$ is injective as the $idx(X_{i})$ are distinct for each $i$ (sim. for $idx(Y_{i})$).
Now we need the following Lemma:
Lemma 3.
For any $\epsilon>0$, and for sufficiently large $d$, there exists at least $M=2^{(\frac{1}{4}-\epsilon)d}$ distinct instantiations of $\{\bm{x}_{i}\}_{0=1}^{M-1}$ and $M$ distinct instantiations of $\{\bm{y}_{i}\}_{i=0}^{M-1}$ of $\bm{Y}$ such that $p_{i}:=\iota(\bm{x_{i}})+\iota(\bm{y_{i}})+1$ are distinct primes, and $\iota(\bm{x_{j}})+\iota(\bm{y_{k}})+1\neq p_{i}$ for any $10\leq i,j,k\leq M-$ 1 except $i=j=k$.
Proof.
We begin by lower bounding the number of prime pairs; that is, the number of instantiations $(\bm{x},\bm{y})$ of $\bm{X},\bm{Y}$ such that $(F(\bm{x},\bm{y}))^{2}=\iota(\bm{x})+\iota(\bm{y})+1$ is prime. Each prime $p$ less than or equal to $2^{d}$ will have exactly 1 prime pair. The number of primes $\pi(m)$ less than or equal to any given integer $m\geq 17$ is lower bounded by $\frac{m}{\ln m}$ (Rosser and Schoenfeld 1962). Thus, we have that the number of prime pairs is at least:
$\frac{2^{d}}{d\ln 2}$ (18)
Given any instantiation $\bm{x}$ of $\bm{X}$, we call $\bm{y}$ a prime completion of $\bm{x}$ if $(\bm{x},\bm{y})$ is a prime pair. We now claim that there are at least $M$ instantiations of $\bm{X}$ such that each has at least $12M^{2}+$ 1 prime completions. Suppose for contradiction this was not the case. Then the total number of prime pairs is upper bounded by:
$(M-1)\times 2^{d-|\bm{X}|}+(2^{|\bm{X}|}-M+1)\times 2M^{2}$
$<2^{(\frac{1}{4}-\epsilon)d}\times 2^{d-|\bm{X}|}+2^{|\bm{X}|}\times 2^{(\frac{1}{2}-2\epsilon)d+1}$
$=2^{(\frac{5}{4}-\epsilon)d-|\bm{X}|}+2^{(\frac{1}{2}-2\epsilon)d+|\bm{X}|+1}$
$\leq 2^{(\frac{11}{12}-\epsilon)d}+2^{(1-2\epsilon)d+1}$
The first line is an upper bound on the number of prime pairs in this case; $(M-1)$ instantiations of $\bm{X}$ with any $\bm{y}$ being a potential prime completion ($2^{d-|\bm{X}|}$ total), and the rest having at most $2M^{2}$ prime completions. The second line follows by substituting $M$, the third by rearrangement, and the fourth using the fact that $\frac{1}{3}d\leq|\bm{X}|\leq\frac{1}{2}d$. But this upper bound is less than the lower bound above, for sufficiently large $d$. Thus, we have a contradiction.
Now, to finish the Lemma, we describe an algorithm for picking the $M$ instantiations $\{\bm{x}_{i}\}_{i=0}^{M-1}$, $\{\bm{y}_{i}\}_{i=0}^{M-1}$. From the claim above, we have $M$ instantiations $\{\bm{x}_{i}\}_{i=0}^{M-1}$ each with at least $12M^{2}+$ 1 prime completions. We iterate over $m=0,...,M-1$. Suppose that at iteration $m$, we have already chosen $\{\bm{y}_{i}\}_{i=0}^{m-1}$ such that $p_{i}:=\iota(\bm{x_{i}})+\iota(\bm{y_{i}})+1$ are distinct primes for $10\leq i\leq m-$ 1, and $\iota(\bm{x_{j}})+\iota(\bm{y_{k}})+1\neq p_{i}$ for any $10\leq j\leq M-$ 1 and $10\leq i,k\leq m-$ 1 except $i=j=k$. For $\bm{x_{m}}$, we aim to choose a prime completion $\bm{y_{m}}$ such that
$(i)\;\iota(\bm{x}_{j})+\iota(\bm{y}_{k})+1\neq\iota(\bm{x}_{m})+\iota(\bm{y}_{m})+1$ (19)
$(ii)\;\iota(\bm{x}_{j})+\iota(\bm{y}_{m})+1\neq p_{k}+1$ (20)
for any $10\leq j\leq M-$ 1 and any $0\leq k\leq m$ except $j=k=m$. Thus, there are at most $2*M*(m+1)\leq 2M^{2}$ values that $\iota(\bm{y}_{m})$ must not take; as we have $12M^{2}+$ 1 prime completions, we can always choose a $\bm{y}_{m}$ satsifying the conditions (i), (ii). Given conditions (i), (ii) together with the inductive hypothesis, we have that $p_{i}:=\iota(\bm{x_{i}})+\iota(\bm{y_{i}})+1$ are distinct primes for $10\leq i\leq(m-1)+$ 1, and $\iota(\bm{x_{j}})+\iota(\bm{y_{k}})+1\neq p_{i}$ for any $10\leq j\leq M-$ 1 and $10\leq i,k\leq(m-1)+$ 1 except $i=j=k$. ∎
With this Lemma in hand, we can finish the argument as follows. By Lemma 3, we have $M$ distinct instantiations $\{\bm{x}_{i}\}_{i=0}^{M-1},\{\bm{y}_{i}\}_{i=0}^{M-1}$ such that $p_{i}:=\iota(\bm{x}_{i})+\iota(\bm{y}_{i})$ is prime for every $i$; suppose that these are ordered such that $p_{0}<...<p_{M-1}$. Now consider the submatrix $\bm{M^{\prime}}\in\mathbb{R}^{M\times M}$ of $\bm{M}_{F(\bm{X},\bm{Y})}$ obtained by taking the rows $(n(\bm{x}_{i}))_{i=0}^{M-1}$ and columns $(n(\bm{y}_{i}))_{i=0}^{M-1}$ (in-order). The rank $rank(\bm{M}_{F(\bm{X},\bm{Y})})$ is lower bounded by $rank(\bm{M}^{\prime})$; thus, we seek to find $rank(\bm{M}^{\prime})$.
Lemma 4.
$rank(\bm{M}^{\prime})=M$
Proof.
This proof is a variation on Example 10 from (Fawzi et al. 2015), but differs from that matrix in that there is no guarantee that the entries are increasing. As such, we rely on a slightly different technique to prove this result.
Recall that $F(\bm{X},\bm{Y})=\pm\sqrt{\iota(\bm{X})+\iota(\bm{Y})+1}$. Thus, the matrix $\bm{M^{\prime}}$ is given by:
$\bm{M}^{\prime}_{ij}=\pm\sqrt{\iota(\bm{x_{i}})+\iota(\bm{y_{j}})+1}$ (21)
Now consider the submatrices $\bm{M^{\prime}}^{(1)},...,\bm{M^{\prime}}^{(M)}$ defined by $\bm{M^{\prime}}^{(i)}:=\bm{M^{\prime}}_{0:i-1,0:i-1}$ (i.e. the first $i$ rows and columns). We show by induction that $\bm{M^{\prime}}^{(i)}$ has rank $i$. The base case $i=1$ is clear.
For the inductive step, suppose that $\bm{M^{\prime}}^{(i-1)}$ has rank $i-1$. Then consider $\bm{M^{\prime}}^{(i)}$. Note that the square of the bottom right entry $(\bm{M^{\prime}}_{i-1,i-1})^{2}=\iota(\bm{x}_{i-1})+\iota(\bm{y}_{i-1})+1=p_{i-1}$ is prime. We now claim that $(\bm{M^{\prime}}_{jk})^{2}$ is not a positive integer multiple of $p_{i}$ for any $10\leq j,k\leq i-$ 1 except $j=k=i-1$.
Firstly, by Lemma 3 there is no $j,k$ such that $(\bm{M^{\prime}}_{jk})^{2}=\iota(\bm{x}_{j})+\iota(\bm{y}_{k})+1=p_{i}$ unless $j=k=i-1$, i.e. a multiple of $1$ is not possible. We further have that:
$\iota(\bm{x}_{j})+\iota(\bm{y}_{k})+1$
$\leq\iota(\bm{x}_{j})+\iota(\bm{y}_{j})+\iota(\bm{y}_{k})+\iota(\bm{x}_{k})+1$
$=p_{j}+p_{k}-1$
$<2p_{i}-1$
Thus $(\bm{M^{\prime}}_{jk})^{2}$ cannot be a positive integer multiple of $p_{i}$.
Now, the determinant of the matrix $\bm{M^{\prime}}^{(i)}$ takes the form $\alpha\bm{M^{\prime}}_{i-1,i-1}+\beta$, where $\alpha$ is the determinant of $\bm{M^{\prime}}^{(i-1)}$, and $\beta$ is a multilinear function of the entries of $\bm{M}^{\prime(i)}$ excluding $\bm{M^{\prime}}_{i-1,i-1}$. Note that both $\alpha$ and $\beta$ are in the extension field $\mathbb{Q}[\sqrt{P_{i}}]$, where $P_{i}$ is defined to be the set of all primes that divide $(\bm{M^{\prime}}_{jk})^{2}$ for some $10\leq j,k\leq i-$ 1 except $j=k=i-1$.
Now, $(\bm{M^{\prime}}_{i-1,i-1})^{2}=p_{i}$ is not in this set $P_{i}$, as we showed that no other $(\bm{M^{\prime}}_{jk})^{2}$ can be a multiple of $p_{i}$. Thus $\bm{M^{\prime}}_{i-1,i-1}=\pm\sqrt{p_{i}}$ is not in this extension field. By the inductive assumption, $\alpha\neq 0$, and so $det(\bm{M^{\prime}}^{(i)})=\alpha\bm{M^{\prime}}_{i-1,i-1}+\beta$ must be nonzero also. Thus $\bm{M}^{\prime(i)}$ has full rank, i.e. rank $i$.
Putting it all together, we have shown that given any square root function $F(\bm{V})=\pm\sqrt{n(\bm{V})+1}$ and any structured-decomposable and smooth circuit $\mathcal{C}^{\prime}$ computing $F$, and any balanced partition $\bm{X},\bm{Y}$ of $\bm{V}$, then for any $\epsilon>0$ and sufficiently large $d$, we have a lower bound $2^{(\frac{1}{4}-\epsilon)d}<rank(\bm{M}^{\prime})<rank(\bm{M}_{F(\bm{X},\bm{Y})})<|\mathcal{C}^{\prime}|^{2}$. ∎
See 1
Proof.
As noted, Inception PCs can express both monotone and squared PCs as special cases. Inception PCs are strictly more expressive efficient than monotone circuits by Theorem 1, as they can express squared PCs; similarly, Inception PCs are strictly more expressive efficient than squared circuits by Theorem 2. ∎
Appendix C Algorithms
Input: Smooth and (structured) decomposable probabilistic circuit with root node $n$
1 if $n$ is input node then
2      return $\texttt{INPUT}(\overline{f_{n}})$
3else if $n$ is product node then
4      return $\texttt{PROD}(\{\texttt{CONJ}(n_{i})\}_{n_{i}\in\textnormal{in}(n)})$
5else if $n$ is sum node then
6      return $\texttt{SUM}(\{\texttt{CONJ}(n_{i})\}_{n_{i}\in\textnormal{in}(n)},\{\overline{\theta_{n,n_{i}}}\}_{n_{i}\in\textnormal{in}(n)})$
Result: Smooth and (structured) decomposable probabilistic circuit with root node $\overline{n}$ such that $f_{\overline{n}}=\overline{f_{n}}$
Algorithm 1 Conjugation $\texttt{CONJ}(n)$
Input: Smooth and (structured) decomposable probabilistic circuit with root node $n$; subset of variables $\bm{W}\subseteq\bm{V}$
1 if $n$ is input node then
2      return $\texttt{INPUT}(\sum_{\bm{w}}f_{n}(\bm{w},\bm{V}\setminus\bm{W}))$
3else if $n$ is product node then
4      return $\texttt{PROD}(\{\texttt{MARG}(n_{i};\bm{W})\}_{n_{i}\in\textnormal{in}(n)})$
5else if $n$ is sum node then
6      return $\texttt{SUM}(\{\texttt{MARG}(n_{i};\bm{W})\}_{n_{i}\in\textnormal{in}(n)},\{\theta_{n,n_{i}}\}_{n_{i}\in\textnormal{in}(n)})$
Result: Smooth and (structured) decomposable probabilistic circuit with root node $n^{\prime}$ such that $f_{n^{\prime}}(\bm{V}\setminus\bm{W})=\overline{f_{n}}$
Algorithm 2 Marginalization $\texttt{MARG}(n;\bm{W})$
Input: Compatible probabilistic circuits with root nodes $n^{(1)},n^{(2)}$
1 if $n^{(1)},n^{(2)}$ are input nodes then
2      return $\texttt{INPUT}(f_{n^{(1)}}\times f_{n^{(2)}})$
3else if $n^{(1)},n^{(2)}$ are product nodes then
4      return $\texttt{PROD}\left(\{\texttt{MULTIPLY}(n^{(1)}_{i},n^{(2)}_{i})\}_{(n^{(1)}_{i},n^{(2)}_{i})\in\texttt{SORT-NODES-BY-SCOPE}(\textnormal{in}(n^{(1)}),\textnormal{in}(n^{(2)}))}\right)$
5else if $n^{(1)},n^{(2)}$ are sum nodes then
6      return $\texttt{SUM}\left(\{\texttt{MULTIPLY}(n^{(1)}_{i},n^{(2)}_{j})\}_{n^{(1)}_{i}\in\textnormal{in}(n^{(1)}),n^{(2)}_{j}\in\textnormal{in}(n^{(2)})},\{\theta_{n^{(1)}_{i}}\theta_{n^{(2)}_{j}}\}_{n^{(1)}_{i}\in\textnormal{in}(n^{(1)}),n^{(2)}_{j}\in\textnormal{in}(n^{(2)})}\right)$
Result: Smooth and structured decomposable probabilistic circuit with root node $n^{\prime}$ such that $f_{n^{\prime}}=f_{n^{(1)}}\times f_{n^{(2)}}$
Algorithm 3 Multiplication $\texttt{MULTIPLY}(n^{(1)},n^{(2)})$
Input: Augmented probabilistic circuit with root node $n$, over variables $\bm{V},\bm{U},\bm{W}$
1 $n^{\prime}\leftarrow\texttt{MARG}(n;\bm{W})$
2 $n^{\prime\prime}\leftarrow\texttt{CONJ}(n^{\prime})$
3 $n^{\prime\prime\prime}\leftarrow\texttt{MULTIPLY}(n^{\prime\prime},n^{\prime})$
4 return $\texttt{MARG}(n^{\prime\prime\prime};\bm{U})$
Result: InceptionPC
Algorithm 4 InceptionPC Construction
For completeness, we provide algorithms for (i) the conjugation operation on PCs with complex weights; and (ii) the procedure of constructing materialized IncPCs. For convenience, we will interchangably refer to circuits $\mathcal{C}$ and their root node $n$. We will write $\texttt{SUM}(C,\Theta)$ to denote the operation of constructing a sum node with inputs $C$ and weights $\Theta$, $\texttt{PROD}(C)$ to denote the operation of constructing a product node with inputs $C$, and $\texttt{INPUT}(f)$ to denote the operation of constructing an input node with function $f$.
In Algorithm 1, we show the algorithm for conjugation, which simply conjugates the weights at sum nodes and input functions. In Algorithm 4, we show the process of constructing materialized IncPCs. This depends on the basic MULTIPLY and MARG on circuits (Vergari et al. 2021), which we reproduce in Algorithms 3, 2. Note that multiplying two circuits while not blowing up the circuit size generally requires them to be compatible (Vergari et al. 2021), [^8] It was recently shown that incompatible circuits can still sometimes be multiplied efficiently (Zhang et al. 2025); that is, compatibility is a sufficient but not necesssary condition. which, intuitively speaking, requires the product nodes in both circuits to decompose their scope in the same way. In Algorithm 2 we multiply a structured decomposable and smooth circuit and its conjugate, which share the same structure and thus are compatible.
Appendix D Log-Sum-Exp Trick with Complex Numbers
To avoid numerical under/overflow, we perform computations in log-space when computing a forward pass of a PC. For complex numbers, this means keeping the modulus of the number in log-space and the argument in linear-space.
Explicitly, given a set of complex numbers $x_{1}=e^{u_{1}+iv_{1}},...,x_{n}=e^{u_{n}+iv_{n}}$ such that the log-modulus $u_{k}\in\mathbb{R}$ and argument $v_{k}\in\mathbb{R}$ are stored in memory, we can compute the log-modulus $u$ and argument $v$ of $x=x_{1}+\ldots+x_{n}$ as follows:
$u$ $=\log(\left|e^{u_{1}-u_{\max}+iv_{1}}+\ldots+e^{u_{n}-u_{\max}+iv_{n}}\right|)+u_{\max}$ (22)
$v$ $=\arg(e^{u_{1}-u_{\max}+iv_{1}}+\ldots+e^{u_{n}-u_{\max}+iv_{n}})$ (23)
where $u_{\max}=\max(u_{1},..,u_{n})$, ||$\cdot$|| ⋅ | is the modulus function for complex numbers, and $\arg$ is the principal value of the argument function (i.e. $\in(-\pi,\pi]$).
Appendix E Experimental Details
For all experiments, we use the Adam optimizer (Kingma and Ba 2015) with learning rate $0.01$. We use a batch size of $64$ for the binary datasets, and $250$ for all image datasets (ImageNet32, ImageNet64). For the binary datasets, we train for 100 epochs and average across 5 runs. For the image datasets, we train for 200 epochs. Model training was performed on NVIDIA A6000 48GB and NVIDIA H100 80GB GPUs.
For the PC input functions, we use categorical inputs for each pixel $V$. For the binary datasets, this corresponds to 2 parameters for each variable $f(V=i)$ for each $i=0,1$, while for the image datasets, for each color channel, we have 256 parameters $f(V=i)$ for each $i=0,...,255$. For monotone PCs, this takes values in $\mathbb{R}^{\geq 0}$, for squared real PCs or real InceptionPCs, this takes values in $\mathbb{R}$, and for squared complex PCs or complex InceptionPCs this takes values in $\mathbb{C}$.
Appendix F Additional Experimental Results for Binary Datasets
We show in Table 4 training negative log-likelihoods on the twenty binary datasets, together with the difference compared with the test negative log-likelihoods shown in Table 2. It can be seen that the complex parameterization of Inception PCs almost always achieves the best training log-likelihoods, sometimes with large margins over the other models. This indicates the added expressive efficiency obtained through the InceptionPC architecture and employing complex rather than non-negative or real parameters. However, they also generally overfit by a larger margin, leading to some cases where another model achieves better test-set performance. Preventing this overfitting, perhaps through regularization, is thus an important consideration when applying these models to small datasets.
Dataset Model
Monotone PC Squared PC Inception PC
Non-negative Real Complex Non-negative Real Complex
nltcs 6.03 (-0.01) 6.02 (-0.00) 6.01 (-0.02) 6.00 (-0.03) 5.96 (-0.04) 6.98 (-0.04) 5.91 (-0.11)
msnbc 6.25 (-0.00) 6.24 (+0.01) 6.11 (+0.01) 6.07(-0.00) 6.06 (+0.01) 6.05 (-0.00) 6.04 (-0.00)
kdd-2k 2.36 (+0.23) 2.35 (+0.25) 2.49 (+0.26) 2.35 (+0.23) 2.34 (+0.22) 2.36 (+0.23) 2.32 (+0.20)
plants 13.44 (-0.26) 13.14 (-0.24) 14.77 (-0.29) 12.80 (-0.33) 12.19 (-0.62) 12.64 (-0.42) 11.81 (-0.95)
jester 52.91 (+0.14) 52.74 (+0.18) 55.00 (+0.49) 52.75 (-0.22) 51.13 (-1.38) 51.13 (-1.47) 49.74 (-3.01)
audio 40.22 (-0.05) 40.04 (-0.04) 41.56 (+0.07) 39.81 (-0.20) 39.08 (-0.80) 39.20 (-0.71) 38.37 (-1.68)
netflix 56.90 (-0.19) 56.69 (-0.16) 57.59 (-0.09) 56.36 (-0.34) 55.21 (-1.31) 55.36 (-1.19) 54.22 (-2.52)
accidents 29.42 (-0.15) 27.76 (-0.17) 28.04 (-0.11) 26.84 (-0.21) 26.01 (-0.69) 27.17 (-0.13) 25.61 (-1.00)
retail 10.85 (-0.14) 10.68 (-0.14) 10.90 (+0.10) 10.80 (-0.15) 10.86 (-0.14) 10.84 (-0.11) 10.78 (-0.17)
pumsb-star 27.97 (-0.01) 24.94 (-0.01) 25.75 (+0.06) 23.94 (+0.04) 23.56 (-0.13) 24.90 (+0.05) 22.80 (-0.23)
dna 79.86 (-0.35) 79.43 (-0.52) 78.80 (-0.35) 77.64 (-0.53) 78.07 (-1.78) 78.00 (-2.11) 75.57 (-4.20)
kosarek 10.90 (+0.13) 10.63 (+0.09) 12.24 (+0.21) 10.65 (+0.06) 10.77 (+0.08) 10.97 (+0.14) 10.56 (-0.04)
msweb 10.34 (-0.10) 9.83 (-0.09) 10.50 (-0.08) 10.05 (-0.12) 10.75 (-0.09) 10.25 (-0.09) 9.96 (-0.14)
book 33.14 (-0.56) 32.62 (-0.70) 36.52 (-0.50) 32.42 (-0.53) 31.81 (-1.70) 33.04 (-1.14) 30.69 (-2.98)
eachmovie 53.75 (+0.92) 52.19 (+0.91) 63.26 (+1.23) 51.14 (-1.19) 49.18 (-1.58) 48.53 (-2.69) 45.29 (-5.12)
web-kb 148.34 (-7.00) 144.82 (-7.02) 155.86 (-6.17) 142.80 (-12.20) 140.60 (-11.14) 143.76 (-9.56) 139.48 (-14.50)
reuters-52 103.90 (+8.68) 100.82 (+8.19) 106.20 (+9.95) 100.80 (+6.90) 99.57 (+6.40) 98.01 (+8.34) 98.22 (+4.42)
20ng 139.72 (-15.33) 137.63 (-15.35) 149.20 (-14.99) 138.17 (-16.62) 136.04 (-18.11) 140.19 (-15.28) 135.42 (-19.76)
bbc 242.90 (-11.08) 239.24 (-11.64) 248.00 (-11.04) 233.90 (-19.23) 232.51 (-18.77) 234.53 (-18.75) 229.84 (-23.53)
ad 16.43 (-0.50) 14.98 (-0.56) 15.69 (-0.63) 14.65 (-0.70) 15.25 (-0.77) 15.25 (-0.56) 14.48 (-0.84)
Table 4: Train negative log-likelihoods on 20 binary datasets, show together with the difference with the test negative log-likelihood (lower is better). A negative difference indicates that the training NLL is less (better) than the test NLL, while a positive difference indicates the opposite.